OpenStudy (anonymous):

Medal Roberto has been studying a type of bacteria that doubles every month. Originally, there were 3 bacterial cells. He wants to know how many there will be after 43 months. Which equation should he use?

OpenStudy (anonymous):

a43 = 3(2)(43) a43 = 3(2)^43 a43 = 3(2)(43 − 1) a43 = 3(2)^43 − 1

OpenStudy (anonymous):

those are the answer choices

OpenStudy (acxbox22):

what do you think?

OpenStudy (anonymous):

first one

OpenStudy (acxbox22):

well you start off with 3 and it DOUBLES every month doubles=2 and there are 43 months

OpenStudy (acxbox22):

how many will there be after 1 month?

OpenStudy (anonymous):

that means its the 3rd choice

OpenStudy (acxbox22):

dont guess..

OpenStudy (anonymous):

its the 2nd choice

OpenStudy (acxbox22):

correct

OpenStudy (acxbox22):

it doubles every month so 2^43

OpenStudy (acxbox22):

and then multiply that by 3 to find how many total bacteria there were

OpenStudy (anonymous):

8.80 e 12

OpenStudy (acxbox22):

you dont need to find that out lol

OpenStudy (anonymous):

oops